OpenStudy (anonymous):

determine weather each pair of equations represents perpendicular lines. show work y= -x+7 y -x =3

OpenStudy (anonymous):

what can you tell me about conditions for two lines to be perpendicular to each other?

OpenStudy (anonymous):

perpendicular means \[m_{1}=-\frac{1}{m_2}\]

OpenStudy (anonymous):

y=mx+c m=-1, m=1 so for perpndicular m1*m2= -1 here the condition satisfied

OpenStudy (anonymous):

find the gradients of both the equations and if the gradient of one equation is the negative reciprocal of the other equations, then they are perpendicular

OpenStudy (anonymous):

first line has slope -1, second has slope 1, so perpendicular.

OpenStudy (anonymous):

equation*